OpenStudy (kitten_is_back):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (anonymous):

If I remember what these questions are about, C, and D are correct. \[-1 \times 1 = -1 \] \[\dfrac{-1}{1} = -1\] Negative one is part of the set. Ergo, closed.

OpenStudy (kitten_is_back):

so times and subtraction?

OpenStudy (anonymous):

Subtraction does not work, since you get 0, and that is not part of the set.

OpenStudy (kitten_is_back):

okay so its a c or d we know its c what about addition?

OpenStudy (anonymous):

Actually we know it is both, C, and D. Any my bad, with addition we get zero, subtraction gives us -2, which is also not in the set.

OpenStudy (kitten_is_back):

okay so its just c d?

OpenStudy (anonymous):

Addition: -1 + 1 = 0. Subtraction: -1 - 1 = -2

OpenStudy (anonymous):

Yes.

OpenStudy (kitten_is_back):

so b c d?

OpenStudy (anonymous):

B gives -2, and the set is only {-1, 1} so no.

OpenStudy (kitten_is_back):

okay M<# thans